se the Laplace transform to solve the given initial-value problem.

y' + 3y = e5t,  y(0) = 2

y(t) =
Answer #1

Consider the initial value problem y + 3y 5t e + yo: 2 + .: We know that aty) = 5Y15) S Y (5) – 410) and h(est) S-5 where T= y(t) = 3 e5t - ē35) + 2ē 38 3E Bes* +(2-1) 8 i ylt) Le5t + 15. ē 34 8 8